OpenStudy (anonymous):

Solve -4x - 5 > -9 and describe the graph of the solution. x > 1; closed circle on 1, shading to the right x < 1; closed circle on 1, shading to the left x < 1; open circle on 1, shading to the left x > 1; open circle on 1, shading to the right

OpenStudy (tkhunny):

What did you do to solve this problem?

OpenStudy (anonymous):

No i need help solving it

OpenStudy (tkhunny):

Right. And the word "help" often implies that you do part of it. Thus my question. What did you do already, so we don't have to cover that material again?

OpenStudy (anonymous):

No Im saying i dont understand how to solve it.

OpenStudy (tkhunny):

Okay, that's not good. Can you solve this for m? 4m = 12

OpenStudy (anonymous):

m=8

OpenStudy (tkhunny):

No, sorry. It requires division. If 4m = 12, then (4m)/4 = 12/4 and m = 3 Let's try another one. 3w + 2 = 11

OpenStudy (anonymous):

oh yeah sorry, i forget.

OpenStudy (anonymous):

w=3

OpenStudy (tkhunny):

Excellent: 3w + 2 = 11 Subtract 2: 3w = 9 Divide by 3: w = 3 One more: 3w + 2 > 11

OpenStudy (anonymous):

3w>9

OpenStudy (anonymous):

oh wait hold on

OpenStudy (anonymous):

w>3

OpenStudy (tkhunny):

Perfect. Now let's tackle your original problem. -4x - 5 > -9 Add 5 -4x > -4 What's next?

OpenStudy (anonymous):

x>-1

OpenStudy (tkhunny):

No, not quite. Two things wrong with it. 1) (-4)/(-4) = +1 2) Division by a Negative Number requires you to do WHAT with the inequality?

OpenStudy (anonymous):

Do i need to flip the sign?

OpenStudy (tkhunny):

-4x > -4 leads by division to x < 1

OpenStudy (anonymous):

so i needed to flip it?

OpenStudy (tkhunny):

I don't know what "flip" means. Let me try to suggest another way that compeltely avoids any confusion. -4x > -4 Add 4 -4x + 4 > 0 Add 4x 4 > 4x Divide by 4 1 > x This way, we never divided by a negative number and no one will have to remember to do anything with the inequality. It just is what it is.

OpenStudy (anonymous):

im still not getting it. but thank you for trying

OpenStudy (tkhunny):

1 > x is the same as x < 1 -- right?

OpenStudy (anonymous):

yes

OpenStudy (tkhunny):

Super. Now, take a good, hard look at the third answer WAY UP at the top.